OpenStudy (amistre64):

spose you got $15000 in an account that pays 8% interest compounded monthly; how much can you withdrawal to make the account last no more than 3 years?

OpenStudy (amistre64):

withdrawal each month to make the account go zero after 3 years is a better wording :)

OpenStudy (tad1):

A = P(!-r)^t

OpenStudy (anonymous):

1 st Month = P (1 + R) - W 2 nd Month = ((P (1 + R) - W) (1 + R)) - W \[= P(1+R)^2-W(1+R)-W\] 3rd Month =\[P(1+R)^3-W(1+R)^2-W(1+R)-W\] \[Nth Month =P(1+R)^N-W\left(1+(1+R)+(1+R)^2+\text{...}.(1+R)^{N-1}\right)\] Using Geometric Sum Formula \[ \left(1+(1+R)+(1+R)^2+\text{...}.(1+R)^{N-1}\right)=\frac{\left(1-(1+R)^N\right)}{1-(1+R)}=-\frac{1-(1+R)^N}{R}\] Putting it back together: \[ P(1+R)^N-W\left(-\frac{1-(1+R)^N}{R}\right)=P(1+R)^N+W\frac{\left(1-(1+R)^N\right) }{R}\] P- Initial amount R- monthly interest rate N= months W= Withdrawal Plugging and Setting the equation to zero and Finding W \[15000\left(1+\frac{.08}{12}\right)^{36}+W\frac{\left(1-\left(1+\frac{.08}{12}\right)^{36}\right) }{\frac{.08}{12}}==0,\] W=470.045

OpenStudy (anonymous):

Excel agrees with my result